OpenStudy (arihangdu):

In triangle ABC, triangle C is a right angle. Find the remaining sides and angles. Round your answer to the nearest tenth. a=3, c=19

OpenStudy (arihangdu):

Use Pythagoras relationship. A^2 + b^2= c^2

OpenStudy (wolf1728):

c would be the hypotenuse so c^2 = a^2 + b^2 19^2 = 3^2 + b^2 361 = 9 + b^2 352 = b^2 b = sq root (352)

OpenStudy (arihangdu):

Ok I found the side c=18.8 but I need to find Angels now

OpenStudy (wolf1728):

yes side c = 18.7616630393 To find the angles you need to use trigonometry

OpenStudy (wolf1728):

It would help to have a graphic - I'll make one

OpenStudy (arihangdu):

Ok

OpenStudy (arihangdu):

I have no idea what to use please help me to solve this

OpenStudy (anonymous):

In triangle ABC, triangle C is a right angle. Find the remaining sides and angles. Round your answer to the nearest tenth. a=3, c=19 sin(A) = 3/19 ------------ B = 90 - A

OpenStudy (anonymous):

ur anwers above

OpenStudy (arihangdu):

What is B== 90-A is ? I lost

OpenStudy (anonymous):

sin(A) = 3/19 ------------ B = 90 - A

OpenStudy (wolf1728):

Okay here's the graphic

OpenStudy (anonymous):

did it at least help you my explanation

OpenStudy (wolf1728):

the sine of angle a = opposite/hypotenuse the sine of angle a = 3/19

OpenStudy (arihangdu):

Ok

OpenStudy (wolf1728):

we need to find what angle has a sine of 0.1578947368 (which is 3/19) To do this we need to find the arc sine of 0.1578947368 Here's a good calculator to use: http://www.1728.org/trigcalc.htm

OpenStudy (arihangdu):

Please help me find that please

OpenStudy (wolf1728):

Okay if you go to the calculator, input 0.1578947368 where it says Input >>> <<<<Input then click "asin" (which means arc sine

OpenStudy (arihangdu):

After that what to solve?

OpenStudy (wolf1728):

You are solving for angle a which equals 9.0847 Degrees Then to solve angle b you calculate 90 - 9.0847 Degrees which equals 80.9153 degrees

OpenStudy (wolf1728):

That should be all the sides and angles you had to solve. Any questions about any of that?

OpenStudy (arihangdu):

What is arcsine means? Is inverse of sine?

OpenStudy (arihangdu):

I'm have only It84

OpenStudy (arihangdu):

Is asine is inverse of sine?

OpenStudy (arihangdu):

I'm still confused on sinA =c/b ................................................ B=90-A. What it means ? Please